OpenStudy (anonymous):

You consider investing in a new hotel. The expected revenues are distributed normally with an average of 3 million dollars and a standard deviation of 1 million dollars. What is the probability that your revenues will be lower than 2 million dollars?

OpenStudy (amistre64):



OpenStudy (amistre64):

since we are 1 sd from the mean this makes life a bit easier

OpenStudy (amistre64):

z = (x-mean)/sd z = (2-3)/1 = -1 do we use a ztable or just do an empirical rule estimate?

OpenStudy (amistre64):

50 - 34 = 20 - 4 = 16% perhaps?

OpenStudy (anonymous):

What you calculated is what I needed (I think). I just do not know how you arrived at 16%...? I think I am supposed to do an empirical rule estimate over here. I am so stupid with such things. I cannot believe it lol

OpenStudy (amistre64):

since the 2 is 1 sd from the mean; that area between 2 and 3 consists of 34% of the data according to the empirical estimates

OpenStudy (amistre64):